OpenStudy (anonymous):

find an for each geometric sequence a1=5,r=3,n=6 and a1=8,r=1/2,n=9

OpenStudy (solomonzelman):

\[a_6=a_1 \times r^5\]

OpenStudy (jdoe0001):

\(\Large \bf a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1}\)

OpenStudy (jdoe0001):

and you're given pretty much all values

OpenStudy (anonymous):

huh which 1 is 1 and 2

OpenStudy (solomonzelman):

\[a_6=5 \times 3^5\]

OpenStudy (anonymous):

ok if that is one what is #2

OpenStudy (jdoe0001):

well, you're given \(\bf a_1\quad r\ and \ \quad n\)

OpenStudy (jdoe0001):

so just plug them in

OpenStudy (jdoe0001):

\(\bf a_1=5,r=3,n=6\qquad a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1}\to a_{\color{red}{ 6}}=5\cdot 3^{{\color{red}{ 6}}-1}\)

OpenStudy (jdoe0001):

for the 2nd one, the same, just plug the values given

OpenStudy (anonymous):

ok tank you so much

OpenStudy (jdoe0001):

yw